OpenStudy (anonymous):

For which equation would x = 12 be a solution? x+ 4 ≤ 12 12 ÷x> 12 12 -x≥ 4 12x< 200

jimthompson5910 (jim_thompson5910):

For each answer choice, replace x with 12 and evaluate. If you get a true inequality, then you found your answer(s) So for instance, in the first answer choice x+4 <= 12 12 + 4 <= 12 ... replace x with 12 16 <= 12 That's FALSE, so x = 12 is NOT a solution to x+4 <= 12

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

so it the third one @Rohangrr

zepdrix (zepdrix):

The third one? 12-x≥4 becomes 12-12≥4 0≥4 Hmm that one appears to be false also doesn't it? :o zero is not greater than or equal to 4.

OpenStudy (anonymous):

sorry i have to go please answer the question i have to go each lunch with my teacher sorry bye

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Keep the work on! all the best

OpenStudy (anonymous):

lol